OpenStudy (anonymous):

Which of the following properties tells us that 5(2x)=(5*2)x ? A. Commutative Property of Multiplication B. Associative Property of Multiplication C. Distributive Property D. None of the above

OpenStudy (zehanz):

It helps if you know what these properties are: Commutative P. of M. is: ab=ba Assoc. P. of M. is a(bc)=(ab)c Distrib. P. of M. is a(b+c)=ab+ac. You cannot miss the right one now... So, apart from choosing the right answer, more important is: Learn the definitions! As soon as you've learned definitions, you'll laugh at this kind of question. They have (almost) nothing to do with math. You just have to know stuff.

OpenStudy (anonymous):

associative law x + (y + z) = (x+y) + z or x(yz) = (xy)z